how do I put y=-x/3 + 8/3 in standard form ? AX + BX = C is the standard form so should it be x/3 + y = 8/3? Can't figure out if I should flip it to 3/x or not?
Answer by Alan3354(69443)

how do I put y=-x/3 + 8/3 in standard form ? AX + BX = C is the standard form so should it be x/3 + y = 8/3? Can't figure out if I should flip it to 3/x or not?
--------------
y=-x/3 + 8/3
I suggest eliminating fractions first.
3y = -x + 8
x + 3y = 8
-------
PS You don't invert, it's still x/3